A body with mass m is acted by two forces F1 and F2 . if m=5.2kg, F1=3.7N, and F2=4.3N, find the vector acceleration of body?

Grade:12th pass

5 Answers

Piyush Kumar Behera
7 years ago
Please specify the direction of the forcess acting on the body.Nor their will no fixed magnitude  nor direction of the acceleration vector.Please do repost the question!!
Amna Saeed
7 years ago
Force f1 acts in upward direction and f2 towards the east ie right side direction. Both forces make an angle of 90 degree between each other. Now that direction are specified plz answer the question
Hamza
7 years ago
F1 is in upward ( North ) and F2 is toward right ( east ). Angle between F1 and F2 is 90°. Now answere this question please... Piyush Kumar.
Piyush Kumar Behera
7 years ago
Ok ,in this question find the net acceleration in the north direction that is F1/m=3.7/5.2=0.711similarly find the net acceleration in the east dithat is that is F2/m=4.3/5.2=0.826.now if you know vectors then as north and east direction are perpendicular so the net magnitude is√(0.711^2+0.826^2)=1.089(approx)this is from the formula of net magnitude of two vectors.Its direction is tan^-1(3.7/4.3) north of east.hope that helps !!please approve this solution to encourage me and if you don`t get the solution then please post it again here.. thanks
inam ur rehman
one year ago
 A body with mass m is acted on by two forces and as
shown in Fig. 4-28. If m -
 5.2 kg, F1 -
 3.7 N, and F2 -
4.3 N, find the vector acceleration of the body.